A sample of nickel metal with a mass of 82.3 grams at 70.0 °C is dropped into 100.0 grams of water at 20.0 °C. If the specific heat of the nickel metal is 0.444 J/g°C and the specific heat of the water is 4.20 J/g°C, then what was the final temperature of the system?


1
Expert's answer
2021-03-10T06:55:52-0500

Solution:

1) This problem can be summarized thusly:

qlost by nickel = qgained by water


q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance

m = mass of sample

C = specific heat capacity (J oC-1 g-1)

Tf = final temperature

Ti = initial temperature


2) Therefore:

qlost by nickel = (82.3 g) × (0.444 J/g°C) × (Tf - 70.0)°C

qlost by nickel = 36.5412 × (Tf - 70.0)


qgained by water = (100.0 g) × (4.20 J/g°C) × (Tf - 20.0)°C

qgained by water = 420 × (Tf - 20.0)


420 × (Tf - 20.0) = 36.5412 × (Tf - 70.0)

11.494 × (Tf - 20.0) = Tf - 70.0

11.494 × Tf - 229.878 = Tf - 70.0

10.494 × Tf = 159.878

Tf = 15.235°C = 15°C


Answer: The final temperature of the system (Tf) is about 15°C.